# Galactic transient sources with the Cherenkov Telescope Array

### v1

@@ Abstract
A wide variety of Galactic sources show transient emission at soft and hard X-ray energies: low-mass and high-mass X-ray binaries containing compact objects (e.g., novae, microquasars, transitional millisecond pulsars, supergiant fast X-ray transients), isolated neutron stars exhibiting extreme variability as magnetars as well as pulsar wind nebulae. Although most of them can show emission up to MeV and/or GeV energies, many have not yet been detected in the TeV domain by Imaging Atmospheric Cherenkov Telescopes. In this paper, we explore the feasibility of detecting new Galactic transients with the Cherenkov Telescope Array (CTA) and the prospects for studying them with Target of Opportunity observations. We show that CTA will likely detect new sources in the TeV regime, such as the massive microquasars in the Cygnus region, low-mass X-ray binaries with low-viewing angle, flaring emission from the Crab pulsar-wind nebula or other novae explosions, among others. We also discuss the multi-wavelength synergies with other instruments and large astronomical facilities.
@@
Timing astronomy and variability studies have proven to be a powerful tool to study extreme astrophysical processes at very high energies (VHE, E 100 GeV). The improvement of the Imaging Atmospheric Cherenkov Technique (IACT) over the past decade has revealed new transient phenomena with variability timescales from seconds to several weeks. The last generation of IACTs have discovered several classes of transient TeV sources such as gamma-ray bursts (GRBs) (MAGIC Collaboration et al., 2019; Abdalla et al., 2019), flaring blazars associated with high-energy neutrino sources (IceCube Collaboration et al., 2018) or Galactic novae (H. E. S. S. Collaboration et al., 2022; Acciari et al., 2022), among others, unveiling new types of VHE emitters with highly variable fluxes.
@@
The Cherenkov Telescope Array (CTA) will be the next generation ground-based observatory for VHE astronomy. It will allow the detection of gamma rays in the 20 GeV-300 TeV domain, with two observatory sites, one in the Northern hemisphere (CTA-N; Observatorio Roque de los Muchachos, La Palma, Spain) and another in the Southern one (CTA-S; Paranal, Chile). It will provide an improved sensitivity with respect to the current generation of IACTs of of about an order of magnitude (Cherenkov Telescope Array Consortium et al., 2019). Of special importance will be the sensitivity of CTA to short-timescale phenomena111For CTA performance, see: https://www.cta-observatory.org/science/ctao-performance/. CTA will have 104-105 better sensitivity than the LAT instrument onboard the Fermi satellite for the detection of short-duration transient events (Funk et al., 2013).
@@
The low energy threshold of 20 GeV of the largest telescopes of the array, the Large-Sized Telescopes (LSTs; CTA-LST Project et al., 2023) is key for the detection of new transient sources at the lower end of the VHE regime. This capability, together with the fast slewing response of the LSTs, which can be re-pointed in about 20 seconds, will allow a swift reaction to transient events. The Medium and Small Sized telescopes (MSTs and SSTs) will also be key to understand the emission of this sources at higher energies. Finally, since the CTA observatory will consist of two arrays located in two hemispheres, it will provide a better and more continuous coverage of many transient events accessible from both sites.
@@
The core program of CTA will consist of different Key Science Projects (KSPs) which were considered to address the science questions of CTA (see Cherenkov Telescope Array Consortium et al., 2019 for more details). The Transients KSP is proposed to encompass the follow-up observations of several classes of targets such as GRBs, gravitational waves (GWs), high-energy neutrinos, core-collapse supernovae (CCSNe) and Galactic transients.
@@
In this paper, we focus on Galactic sources hosting compact objects whose emission is not periodic and/or that display unexpected flaring events, outflows or jets as described in Cherenkov Telescope Array Consortium et al. (2019). Extragalactic transient events such as GRBs, core-collapse SNe or GWs will be addressed in separate publications. We discuss the capabilities of CTA to detect new transient phenomena at VHE from sources of Galactic origin, ranging from microquasars, to pulsar-wind nebulae (PWNe) flares, to novae, transitional millisecond pulsars or magnetars among others. For our simulations, we have used the software packages ctools
@@
222ctools+ is a software specifically developed for the scientific analysis of gamma-ray data, see http://cta.irap.omp.eu/ctools/index.html (Knödlseder et al., 2016) and
@@
333 Gammapy— is an open-source Python package developed for gamma-ray astronomy, see https://gammapy.org/ (Donath et al., 2023; Aguasca-Cabot et al., 2023) with the official CTA observatory instrument response functions (IRFs).444The IRF version
@@
prod5+ version corresponding to the Alpha Configuration, which corresponds to the first stage of CTA observatory construction has also been tested in some science cases and are specified in the text. For a full description of CTA observatory IRFs and configurations see Maier et al. (2023).
The sources of our interest are described in the following Subsections 1.1-1.6. We present the sensitivity of CTA to Galactic transient detection in Section 2 and population studies in Section 3. The simulations, analysis results and discussion for each type of transient are collected in Section 4. Section 5 describes the synergies with multi-wavelength and multi-messenger astronomical facilities. The summary and final conclusions are listed in Section 6.
@@ 1.1 Microquasars
Up to now, only three microquasars have been detected in the high-energy (HE, E100 MeV) domain: Cygnus X-1 (Bulgarelli et al., 2010; Sabatini et al., 2010, 2013; Malyshev et al., 2013; Zanin et al., 2016; Zdziarski et al., 2017), Cygnus X-3 (Tavani et al., 2009; Fermi LAT Collaboration et al., 2009; Zdziarski et al., 2018), and SS 433 (Bordas et al., 2015; Sun et al., 2019; Rasul et al., 2019; Xing et al., 2019; Li et al., 2020). Steady VHE emission was detected from the interaction regions between the jet and the surrounding nebula in SS 433 (Abeysekara et al., 2018a; H. E. S. S. Collaboration et al., 2024). Furthermore, HAWC has announced the TeV detection of the microblazar V4641 Sgr with 7 years of data (Tibolla, 2023). A strong hint of transient emission was found at VHE in Cygnus X-1 (Albert et al., 2007). Strong hints of HE emission were also found in the low-mass X-ray binary V404 Cyg during its 2015 outburst (Loh et al., 2016; Piano et al., 2017). The expectations for the detection of both massive microquasars and low-mass X-ray binaries (LMXBs) with CTA are presented in Sections 4.1-4.2.
The relevance for studying gamma-ray binaries and microquasars has already been addressed by Paredes et al. (2013). In this paper, we do not focus on gamma-ray binaries displaying periodic orbital variability, but only on microquasars, to better investigate the potential VHE emission of this class of binaries for which only SS 433 has been confirmed as a VHE emitter. We discuss high-mass microquasars (Section 4.1) separately to low-mass X-ray binaries (Section 4.2).
@@ 1.2 Transitional millisecond pulsars
Transitional millisecond pulsars (tMSPs) are a class of neutron star binaries that has emerged in the last decade with the discoveries of three confirmed systems: PSR J1023+0038 (Archibald et al., 2009a; Patruno et al., 2014a), XSS J1227-4853 (de Martino et al., 2010; Bassa et al., 2014) and IGR J1824-2452 in the globular cluster M28 (Papitto et al., 2013). Additionally, a handful of candidate tMSPs have been recently discovered in the X-ray and GeV ranges (see review by Papitto & de Martino, 2022). tMSPs alternate between a radio-loud MSP state (RMSP, showing radio pulsations and no sign of an accretion disk) and a sub-luminous LMXB state (forming an accretion disk and showing X-ray pulsations). These sources are the direct link between the LMXB and the radio MSP phases in which neutron stars are spun up to ms periods during the LMXB-phase. Sudden transitions between the two states occur on a timescale of a few days to weeks, and are accompanied by drastic changes across the electromagnetic spectrum. The transition from the RMSP to LMXB state is accompanied by brightening of optical, UV (Patruno et al., 2014b; Takata et al., 2014), X-ray and gamma-ray (Stappers et al., 2013) emission with the disappearance of radio pulsations. The origin of these transitions is still debated and, for this, intense multi-wavelength campaigns are on-going to understand the phenomenology in both the RMSP and LMXB states. tMSPs were so far not detected in the VHE regime. The upper limits (ULs) in the VHE regime from tMSPs are discussed in Section 4.3.
@@ 1.3 Pulsar wind nebulae
Pulsar wind nebulae (PWNe) are bubbles or diffuse structures of relativistic plasma powered by a central highly-magnetized rotating neutron star (pulsar). PWNe show three different evolution stages (see e.g. Olmi, 2023 and references therein): a free-free expansion stage, in which there is no interaction with the surrounding supernova remnant (SNR); a reverberation phase in which the supernova reverse shock interacts with the PWN and a post-reverberation or bow shock stage, in which the pulsar leaves the SNR onto the external medium, creating a cometary-like structure. Some PWNe have been found to display TeV halos (Abeysekara et al., 2017; Cao et al., 2023a), which are suggested to be the result of particle escape in evolved pulsar. Recently, several PWNe have been suggested to be PeV particle (leptons) accelerators, with the detection of gamma rays at E100 TeV (Cao et al., 2021, 2023a). The Crab nebula is the standard candle at VHE and both the nebula and the pulsar have been intensively studied. Pulsations have been measured up to TeV energies (Ansoldi et al., 2016) and the nebula spectrum has been detected up to 100 TeV by IACTs (MAGIC Collaboration et al., 2020) and recently extended to PeV (Lhaaso Collaboration et al., 2021). Unexpectedly, the Crab nebula displays rapid flaring emission over daily timescales at HE as reported by AGILE and Fermi-LAT (Tavani et al., 2011; Abdo et al., 2011). The enhanced fluxes measured over different flaring episodes were a factor three-to-six times larger than the standard flux. So far, no signs of variability have been reported at VHE (Mariotti, 2010; Ong, 2010; H. E. S. S. Collaboration et al., 2014; van Scherpenberg et al., 2019). The characterization of the expected VHE emission to be putatively detected by CTA is shown in Section 4.4.
@@ 1.4 Novae
Novae are thermonuclear runaway explosions on the surface of a white dwarf star in binary systems involving a white dwarf accreting matter, often through an accretion disc, usually from a late-type star (Gallagher & Starrfield, 1978). They are detected as transient events exhibiting huge and sudden increase of brightness. Though novae have been studied both observationally and theoretically for many decades, a comprehensive understanding of nova physics is still lacking (Iben, 1982; Yaron et al., 2005). Particle acceleration in novae was predicted before the launch of the Fermi Gamma-ray space telescope (see Tatischeff & Hernanz, 2007). Shortly after, GeV emission from the outburst of the symbiotic binary system V407 Cygni, comprised of a white dwarf and an evolved red giant companion, was first detected. Subsequently, classical novae with main sequence donor stars were also detected (Abdo et al., 2010a; Ackermann et al., 2014).666Gomez-Gomar et al. (1998) (and references therein) predicted gamma-ray emission from novae but of nuclear origin, in the keV-MeV domain. More recently, VHE emission in novae has been predicted and searched for in a handful of sources (see e.g. Aliu et al., 2012; Ahnen et al., 2015), with the first detection at VHE gamma rays occurring in 2021 in the recurrent nova RS Ophiuchi (RS Oph, Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022; Abe et al., 2023).
@@
Since the first detection at HE gamma rays from nova Cygni 2010, 19 novae777According to https://asd.gsfc.nasa.gov/Koji.Mukai/novae/latnovae.html (as of February 2024). have been detected in this energy band (only RS Oph at VHE) with a rate of about one outburst detection per year. All novae so far detected at HE have been bright in the visible band (), and the vast majority are nearby sources with distances within (Franckowiak et al., 2018). Non-thermal emission is expected to arise from leptonic and hadronic interactions by particles accelerated in radiative expanding shocks (Abdo et al., 2010a; Hernanz & Tatischeff, 2012), which can originate from the interaction of the ejecta during the initial stage of the outburst and the circumbinary material, or with the fast wind produced by nuclear burning in later stages of the outburst (Abdo et al., 2010a; Ackermann et al., 2014; Martin et al., 2018).
@@
Based on observations of novae in the nearby M31 Galaxy, as well as binary population synthesis models for the Milky Way, a rate of approximately 30 nova events per year is expected (see Section 4.5). However, a significant proportion of these will be obscured by intervening dust in the Galactic plane. The number of nova events that will be detectable at HE and VHE gamma rays will be further constrained by properties of the system, such as the shock velocity and the target material density. This dependence on the parameter space and prospects for detection of novae at VHE will be characterised for CTA in Section 4.5.

### v2

@@ Abstract
A wide variety of Galactic sources show transient emission at soft and hard X-ray energies: low-mass and high-mass X-ray binaries containing compact objects, isolated neutron stars exhibiting extreme variability as magnetars as well as pulsar wind nebulae. Although most of them can show emission up to MeV and/or GeV energies, many have not yet been detected in the TeV domain by Imaging Atmospheric Cherenkov Telescopes. In this paper, we explore the feasibility of detecting new Galactic transients with the Cherenkov Telescope Array Observatory (CTAO) and the prospects for studying them with Target of Opportunity observations. We show that CTAO will likely detect new sources in the TeV regime, such as the massive microquasars in the Cygnus region, low-mass X-ray binaries with low-viewing angle, flaring emission from the Crab pulsar-wind nebula or other novae explosions, among others. Since some of these sources could also exhibit emission at larger timescales, we additionally test their detectability at longer exposures. We finally discuss the multi-wavelength synergies with other instruments and large astronomical facilities.
@@
Timing astronomy and variability studies have proven to be a powerful tool to study extreme astrophysical processes at very high energies (VHE, E 100 GeV). The improvement of the Imaging Atmospheric Cherenkov Technique (IACT) over the past decade has revealed new transient phenomena with variability timescales from seconds to several weeks. The last generation of IACTs have discovered several classes of transient TeV sources such as gamma-ray bursts (GRBs) (MAGIC Collaboration et al., 2019; Abdalla et al., 2019), flaring blazars associated with high-energy neutrino sources (IceCube Collaboration et al., 2018) or Galactic novae (H. E. S. S. Collaboration et al., 2022; Acciari et al., 2022), among others, unveiling new types of VHE emitters with highly variable fluxes (Carosi & López-Oramas, 2024).
@@
The Cherenkov Telescope Array Observatory (CTAO) will be the next generation ground-based observatory for VHE astronomy. It will allow the detection of gamma rays in the 20 GeV-300 TeV domain, with two observatory sites, one in the Northern hemisphere (CTAO-N; Observatorio Roque de los Muchachos, La Palma, Spain) and another in the Southern one (CTAO-S; Paranal, Chile). It will provide an improved sensitivity with respect to the current generation of IACTs of of about an order of magnitude (Cherenkov Telescope Array Consortium et al., 2019). Of special importance will be the sensitivity of CTAO to short-timescale phenomena111For CTAO performance, see: https://www.CTAO.org/for-scientists/performance/.
CTAO will have 104-105 better sensitivity than the LAT instrument onboard the Fermi satellite for the detection of short-duration transient events (Funk et al., 2013).
@@
The low energy threshold of 20 GeV of the largest telescopes of the array, the Large-Sized Telescopes (LSTs; CTA-LST Project et al., 2023) is key for the detection of new transient sources at the lower end of the VHE regime. This capability, together with the fast slewing response of the LSTs, which can be re-pointed in about 20 seconds, will allow a swift reaction to transient events. The Medium and Small Sized telescopes (MSTs and SSTs) will also be key to understand the emission of this sources at higher energies. Finally, since the CTAO observatory will consist of two arrays located in two hemispheres, it will provide a better and more continuous coverage of many transient events accessible from both sites.
@@
The core program of CTAO will consist of different Key Science Projects (KSPs) which were considered to address the science questions of CTAO (see Cherenkov Telescope Array Consortium et al., 2019 for more details). The Transients KSP is proposed to encompass the follow-up observations of several classes of targets such as GRBs, gravitational waves (GWs), high-energy neutrinos, core-collapse supernovae (CCSNe) and Galactic transients.
@@
In this paper, we focus on Galactic sources hosting compact objects whose emission is not periodic and/or that display unexpected flaring events, outflows or jets as described in the Galactic transients section of the Transients KSP as defined in Cherenkov Telescope Array Consortium et al. (2019). Extragalactic transient events such as GRBs, core-collapse SNe or GWs will be addressed in separate publications. We discuss the capabilities of CTAO to detect new transient phenomena at VHE from sources of Galactic origin, ranging from microquasars, to pulsar-wind nebulae (PWNe) flares, to novae, transitional millisecond pulsars or magnetars among others. Some of these sources could also exhibit persistent emission, hence we additionally test the detectability at longer exposures in some specific cases. Since the nature of the source classes of study and hence the physical processes are different, the simulated timescales of the expected VHE emission also vary. We have used timescales ranging from as low as 10 min to few hours for transient detection and up to 50-200 h to test persistent emission in certain sources of interest. For our simulations, we have used the software packages ctools
@@
222ctools+ is a software specifically developed for the scientific analysis of gamma-ray data, see http://CTAO.irap.omp.eu/ctools/index.html (Knödlseder et al., 2016) and
@@
333 Gammapy— is an open-source Python package developed for gamma-ray astronomy, see https://gammapy.org/ (Donath et al., 2023; Aguasca-Cabot et al., 2023) with the official CTAO observatory instrument response functions (IRFs).444The IRF version
@@
prod5+ version corresponding to the Alpha Configuration, which corresponds to the first stage of CTAO observatory construction has also been tested in some science cases and are specified in the text. For a full description of CTAO observatory IRFs and configurations see Maier et al. (2023).
The source classes of our interest are described in the following Subsections 1.1-1.5. We present the sensitivity of CTAO to Galactic transient detection in Section 2 and population studies in Section 3. The simulations, analysis results and discussion for each type of transient are collected in Section 4. Section 5 describes the synergies with multi-wavelength and multi-messenger astronomical facilities. The summary and final conclusions are listed in Section 6.
@@ 1.1 Microquasars
Up to now, only three microquasars have been detected in the high-energy (HE, E100 MeV) domain: Cygnus X-1 (Bulgarelli et al., 2010; Sabatini et al., 2010, 2013; Malyshev et al., 2013; Zanin et al., 2016; Zdziarski et al., 2017), Cygnus X-3 (Tavani et al., 2009; Fermi LAT Collaboration et al., 2009; Zdziarski et al., 2018), and SS 433 (Bordas et al., 2015; Sun et al., 2019; Rasul et al., 2019; Xing et al., 2019; Li et al., 2020), all of them hosting a massive companion star. In the case of low-mass microquasars, the only one that has displayed a strong hint of gamma-ray emission (at high energies) was the binary V404 Cyg during its 2015 outburst (Loh et al., 2016; Piano et al., 2017). Steady VHE emission was first detected from the interaction regions between the jet and the surrounding nebula for the time in SS 433 (Abeysekara et al., 2018a; H. E. S. S. Collaboration et al., 2024). Other microquasars have been recently reported to be sources of persistent of TeV and PeV emission (for details see i.e. Alfaro et al., 2024; LHAASO Collaboration, 2024). Regarding flaring emission, the strongest hint reported was the 4.1 transient signal (post-trial) found at VHE in Cygnus X-1 (Albert et al., 2007).The expectations for the detection of both massive microquasars and low-mass X-ray binaries (LMXBs) with CTAO are presented in Sections 4.1-4.2.
The relevance for studying binary systems in the VHE regime has already been addressed by Paredes et al. (2013); Chernyakova et al. (2019). In this paper, we do not focus on gamma-ray binaries displaying periodic orbital variability and likely powered by non-accreting pulsars, but only on systems powered by accretion and displaying jets, to better investigate the potential VHE emission of this specific class of binaries.We discuss high-mass microquasars (Section 4.1) separately to low-mass X-ray binaries (Section 4.2).
@@ 1.2 Transitional millisecond pulsars
Transitional millisecond pulsars (tMSPs) are a class of neutron star binaries that has emerged in the last decade with the discoveries of three confirmed systems: PSR J1023+0038 (Archibald et al., 2009a; Patruno et al., 2014a), XSS J1227-4853 (de Martino et al., 2010; Bassa et al., 2014) and IGR J1824-2452 in the globular cluster M28 (Papitto et al., 2013). Additionally, a handful of candidate tMSPs have been recently discovered in the X-ray and GeV ranges (see review by Papitto & de Martino, 2022). tMSPs alternate between a radio-loud MSP state (RMSP, showing radio pulsations and no sign of an accretion disk) and a sub-luminous LMXB state (forming an accretion disk and showing X-ray pulsations). These sources are the direct link between the LMXB and the radio MSP phases in which neutron stars are spun up to ms periods during the LMXB-phase. Sudden transitions between the two states occur on a timescale of a few days to weeks, and are accompanied by drastic changes across the electromagnetic spectrum. The transition from the RMSP to LMXB state is accompanied by brightening of optical, UV (Patruno et al., 2014b; Takata et al., 2014), X-ray and gamma-ray (Stappers et al., 2014) emission with the disappearance of radio pulsations. The origin of these transitions is still debated and, for this, intense multi-wavelength campaigns are on-going to understand the phenomenology in both the RMSP and LMXB states. tMSPs were so far not detected in the VHE regime. The constrains to the VHE emission from tMSPs during the LMXB state are discussed in Section 4.3.
@@ 1.3 Pulsar wind nebulae
Pulsar wind nebulae (PWNe) are bubbles or diffuse structures of relativistic plasma powered by a central highly-magnetized rotating neutron star. They represent one of the largest Galactic populations at VHE. Recently, several PWNe have been suggested to be PeV particle (leptons) accelerators, with the detection of gamma rays at E100 TeV (Cao et al., 2021, 2023). The Crab nebula is the standard candle at VHE and both the nebula and the pulsar have been intensively studied. Pulsations have been measured up to TeV energies (Ansoldi et al., 2016) and the nebula spectrum has been detected up to 100 TeV by IACTs (MAGIC Collaboration et al., 2020) and recently extended to PeV (Lhaaso Collaboration et al., 2021). Unexpectedly, the Crab nebula displays rapid flaring emission over daily timescales at HE as reported by AGILE and Fermi-LAT (Tavani et al., 2011; Abdo et al., 2011). The enhanced fluxes measured over different flaring episodes were a factor three-to-six times larger than the standard flux. These episodes of enhanced HE emission have been detected up to 10 GeV (as reported by Tavani et al., 2011) and can last up to few weeks. A detection of the synchrotron tail at higher energies or a additional inverse Compont component in the TeV domain could be expected. So far, no signs of variability have been reported at VHE (Mariotti, 2010; Ong, 2010; H. E. S. S. Collaboration et al., 2014; van Scherpenberg et al., 2019). The characterization of the expected VHE emission to be putatively detected by CTAO is shown in Section 4.4.
@@ 1.4 Novae
Novae are thermonuclear runaway explosions on the surface of a white dwarf star in binary systems involving a white dwarf accreting matter, often through an accretion disc, usually from a late-type star (Gallagher & Starrfield, 1978). They are detected as transient events exhibiting huge and sudden increase of brightness. Though novae have been studied both observationally and theoretically for many decades, a comprehensive understanding of nova physics is still lacking (Iben, 1982; Yaron et al., 2005; Bode & Evans, 2008; Kato et al., 2014; Chomiuk et al., 2021). Particle acceleration in novae was predicted before the launch of the Fermi Gamma-ray space telescope (see Tatischeff & Hernanz, 2007). Shortly after, GeV emission from the outburst of the symbiotic binary system V407 Cygni, comprised of a white dwarf and an evolved red giant companion, was first detected. Subsequently, classical novae with main sequence donor stars were also detected (Abdo et al., 2010a; Ackermann et al., 2014).666Gomez-Gomar et al. (1998) (and references therein) predicted gamma-ray emission from novae but of nuclear origin, in the keV-MeV domain. More recently, VHE emission in novae has been predicted and searched for in a handful of sources (see e.g. Aliu et al., 2012; Ahnen et al., 2015), with the first detection at VHE gamma rays occurring in 2021 in the recurrent nova RS Ophiuchi (RS Oph, Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022; Abe et al., 2025).
@@
Since the first detection at HE gamma rays from nova Cygni 2010, 19 novae777According to https://asd.gsfc.nasa.gov/Koji.Mukai/novae/latnovae.html (as of February 2024). have been detected in this energy band (only RS Oph at VHE) with a rate of about one outburst detection per year. All novae so far detected at HE have been bright in the visible band (), and the vast majority are nearby sources with distances within (Franckowiak et al., 2018). Non-thermal emission is expected to arise from leptonic and hadronic interactions by particles accelerated in radiative expanding shocks (Abdo et al., 2010a; Hernanz & Tatischeff, 2012), which can originate from the interaction of the ejecta during the initial stage of the outburst and the circumbinary material, or with the fast wind produced by nuclear burning in later stages of the outburst (Abdo et al., 2010a; Ackermann et al., 2014; Martin et al., 2018). The VHE signal reported by Acciari et al., 2022; H. E. S. S. Collaboration et al., 2022; Abe et al., 2025 is suggested to be of hadronic origin, due to protons accelerated in the nova shock.
@@
Based on observations of novae in the nearby M31 Galaxy, as well as binary population synthesis models for the Milky Way, a rate of approximately 30 nova events per year is expected (see Section 4.5). However, a significant proportion of these will be obscured by intervening dust in the Galactic plane, preventing multi-wavelength follow-up observations. The number of nova events that will be detectable at HE and VHE gamma rays will be further constrained by properties of the system, such as the shock velocity and the target material density. This dependence on the parameter space and prospects for detection of novae at VHE will be characterised for CTAO in Section 4.5.